Ling invested some money at 7% interest. Ling also invested $74 more than 4 times that amount at 13%. How much is invested at each rate if Ling receives $828.54 in interest after one year? (Round to two decimal places if necessary)
Answer by mananth(16949)

Ling invested some money at 7% interest. Ling also invested $74 more than 4 times that amount at 13%. How much is invested at each rate if Ling receives $828.54 in interest after one year? (Round to two decimal places if necessary)
7% interest ------------- x
13% interest-------------$74+4x
$828.54 interest after one year
7x+13(4x+74)=828.54*100
7x+52x+962=82854
59x=82854-962
59x=81892
/59
x= $1,388.00 ----- 7%
4x+74 = $5626-----13%